250 grams of silver nitrate react with an excess amount of potassium chloride to produce potassium nitrate and silver chloride. How many grams of silver chloride do you get?
1
Expert's answer
2020-05-03T15:11:43-0400

Let's start by writing the balanced reaction equation:

AgNO3 + KCl "\rightarrow" AgCl"\downarrow" + KNO3.

As you can see, when one mole of AgNO3 reacts, 1 mole of silver chloride is produced:

"n(AgNO_3) = n(AgCl)" .

The number of the moles is the ratio of the mass "m" to the molar mass "M". The molar mass of AgNO3 is 107.87+14.01+16.00·3=169.87 g/mol. The molar mass of AgCl is: 107.87+35.45=143.32 g/mol. The mass of silver chloride can be expressed as:

"m(AgCl) = n(AgCl)\u00b7M(AgCl)"

"m(AgCl) = n(AgNO_3)\u00b7M(AgCl)"

"m(AgCl) = \frac{m(AgNO_3)}{M(AgNO_3)}·M(AgCl)"

"m(AgCl) = \frac{250}{169.87}·143.32 = 210.9" g.

Answer: When 250 grams of silver nitrate react with an excess amount of potassium chloride to produce potassium nitrate and silver chloride, you get 210.9 g of silver chloride.